### v1

Abstract
Let $R=k[x_{1},\ldots,x_{d}]$ denote a standard graded polynomial ring over an algebraically closed field $k$, and let $I\subset R$ be a perfect ideal of codimension $2$ with an $n\times(n-1)$ linear presentation matrix $\varphi$. We prove an extended formulation of a theorem of Morey and Ulrich in the case where $I$ satisfies condition $G_{d-1}$, but not condition $G_{d}$.
1 Statement of the problem
Let $R=k[x_{1},\ldots,x_{d}]$ denote a standard graded polynomial ring over an infinite field $k$, and let $I\subset R$ be a perfect ideal of codimension $2$ with an $n\times(n-1)$ linear presentation matrix $\varphi$ – heretofore informally called the associated Hilbert–Burch matrix. Thus, $I$ is generated by the $n$ maximal minors of $\varphi$.
Then, there is a unique $d\times(n-1)$-matrix $B$, whose entries are linear forms in $T=k[{\bf t}]=k[t_{1},\ldots,t_{n}]$, such that
${\bf t}\cdot\varphi={\bf x}\cdot B.$ (1)
$B$ is often called the Jacobian dual matrix of $\varphi$.
S. Morey and B. Ulrich have proved the following beautiful theorem:
Theorem 1.1.
([11, Theorem 1.3]) Let $R$ and $I\subset R=k[x_{1},\ldots,x_{d}]$ be as above, and such that $\mu(I)>d$. If $I$ satisfies condition $G_{d}$, then::::
  1. (i)
    The analytic spread $\ell(I)$ is maximal (i.e., $\ell(I)=d$).
  2. (ii)
    The minimal reduction number of $I$ has the “expected” value (i.e., $\ell(I)-1$).
  3. (iii)
    The Rees algebra $\mathcal{R}_{R}(I)$ of $I$ is Cohen–Macaulay with homogeneous defining ideal of the “expected” form $\langle I_{1}({\bf t}\cdot\varphi),I_{d}(B)\rangle$.
The main goal of this paper is to weaken the crucial hypothesis of this theorem to see what is left in one piece. For this we bring up the following notion introduced in [5]. Let $R$ be a standard graded ring over a field $k$, with irrelevant maximal ideal $\mathfrak{m}$, and $I\subset R$ an equigenerated ideal thereof. Then the special fiber $\mathcal{F}_{R}(I)=\mathcal{R}_{R}(I)/\mathfrak{m}\mathcal{R}_{R}(I)$ can be identified with the $k$-subalgebra $k[It]\subset R[It]$. In this way, the homogeneous defining ideal $Q$ of the latter is contained in a defining ideal of the Rees algebra $\mathcal{R}_{R}(I)$. One then says that the ideal $I$ (or $\mathcal{R}_{R}(I)$) is of fiber type if a defining ideal of $\mathcal{R}_{R}(I)$ has the form $\langle I_{1}({\bf t}\cdot\varphi),Q\rangle$.
In this regard, assuming throughout that $k$ is algebraically closed, we prove:
Theorem 1.2.
Let $R$ and $I\subset R=k[x_{1},\ldots,x_{d}]$ be as above, and such that $\mu(I)>d$ and $I_{1}(\varphi)=\langle x_{1},\ldots,x_{d}\rangle$. If $I$ satisfies condition $G_{d-1}$, but not condition $G_{d}$, then::::
  1. (a)
    The analytic spread $\ell(I)$ is maximal (i.e., $\ell(I)=d$).
  2. (b)
    The Rees algebra $\mathcal{R}_{R}(I)$ of $I$ is of fiber type, but not of the “expected” form (in particular, the inclusion $\langle I_{1}({\bf t}\cdot\varphi),I_{d}(B)\rangle\subset\langle I_{1}({\bf t}\cdot\varphi),Q\rangle$ is proper).
Concerning item (a) we will prove, more strongly, the following generalization of [4, Theorem 2.4 (a)]:
Proposition 1.3.
With the same notation as above, let $I_{n-1}$ denote the linear span of $I$. Then the rational map $\operatorname{\mathbb{P}}^{d-1}\dasharrow\operatorname{\mathbb{P}}^{n-1}$ defined by $I_{n-1}$ is birational onto the image.
As regards the Cohen–Macaulay property of the Rees algebra in item (b), it may fail quite commonly even for $d=3$, as noticed in [4, Remark 4.14]. Since the notion of an expected reduction number is related to this property, it is no clear what would be a replacement for such a notion over here. In any case, it will be clear that the reduction number is bounded below by $d-2$, and there are examples where this number is $d-1$.
Particular cases of Theorem 1.2 have been dealt with very recently ([7]), while some other aspects of the present work have been considered in [2]. Of course, much of what we learned was inspired by earlier work, such as [9], [10], [4].
2 Notation and main tools
Let $R$ denote a Noetherian ring and let $E$ be a finitely generated $R$-module having rank. We give a brief review of certain graded algebras associated with the pair $(R,E)$. For further details, the reader is referred to [15].
  1. $\bullet$
    The symmetric algebra $\mathcal{S}_{R}(E)=\bigoplus_{i\geq 0}\mathcal{S}_{R,i}(E)$, as the direct sum of the symmetric powers of $E.$
  2. $\bullet$
    The Rees algebra $\mathcal{R}_{R}(E)=\mathcal{S}_{R}(E)/\langle R{\rm-torsion}\rangle$.
    This notion of Rees algebra of a module is kosher enough whenever $E$ has a rank, and moreover boils down to the usual definition in the case of an ideal $I=\langle f_{1},\ldots,f_{n}\rangle$: $\mathcal{R}_{R}(I)=\bigoplus_{i\geq 0}I^{i}\simeq R[f_{1}t,\ldots,f_{n}t]\subset R[t]$, as the direct sum of the powers of $I.$
    If, moreover, $R$ is local with maximal ideal $\mathfrak{m}$ (or standard graded with maximal irrelevant ideal $\mathfrak{m}$), we set:
  3. $\bullet$
    The special fiber (or fiber cone) $\mathcal{F}_{R}(E):=\mathcal{R}_{R}(E)/\mathfrak{m}\mathcal{R}_{R}(E).$
In order to represent $\mathcal{S}_{R}(E)$ as a residue algebra of a polynomial ring, we pick a (finite) set of, say, $m$ generators of $E$ over $R$, and the corresponding finite free presentation of $E$ over $R$
$R^{n}\stackrel{{\scriptstyle\varphi}}{{\longrightarrow}}R^{m}\longrightarrow E\rightarrow 0.$ (2)
Then the universal property of the symmetric algebra implies the exact sequence of symmetric algebras
$\mathcal{S}_{R}(R^{n})\stackrel{{\scriptstyle\mathcal{S}_{R}\varphi}}{{\longrightarrow}}\mathcal{S}_{R}(R^{m})\longrightarrow\mathcal{S}_{R}(E)\rightarrow 0,$
itself giving an isomorphism of $R$-modules $\mathcal{S}_{R}(E)\simeq R[{\bf t}]/\langle I_{1}({\bf t}\cdot\varphi)\rangle$, where $R[{\bf t}]$ is a polynomial ring representing $\mathcal{S}_{R}(R^{m})$.
A similar representation of the Rees algebra right off the definition. Namely, drawing upon the above representation of $\mathcal{S}_{R}(E)$, we note that its $R{\rm-torsion}$ is given by
$\tau(E):=\bigcup_{g\in R\setminus\mathcal{Z}(R)}I_{1}({\bf t}\cdot\varphi)\colon g,$
where $\mathcal{Z}(R)$ is the set of sero-divisors of $R$. Thus, $\mathcal{R}_{R}(E)\simeq R[{\bf t}]/\langle I_{1}({\bf t}\cdot\varphi),\tau(E)\rangle$.
In general the complementary generators afforded by $\tau(E)$ can be quite involved even if $E$ is an ideal. If $R$ is a standard graded ring over a field $k$ and $I\subset R$ happens to be an equigenerated ideal thereof, then the special fiber $\mathcal{F}_{R}(I)$ can be identified to the $k$-subalgebra $k[It]\subset R[It]$. In this way, the homogeneous defining ideal $Q$ of the latter is contained in $\tau(I)$.
There is a couple of ways to define the condition $G_{s}$, originally introduced in [1] in the case of ideals. Thus, let $R$ denote a Noetherian ring and let $E$ be a finitely generated $R$-module having rank $e$. We will adopt any one of the following notions.
Definition 2.1.
Given an integer $s\geq 1$, $E$ satisfies condition $G_{s}$ if $\mu(E_{\mathfrak{p}})\leq\dim R_{\mathfrak{p}}+e-1$ for every ${\mathfrak{p}}\in\operatorname{Spec}(R)$ with $11\leq\dim R_{\mathfrak{p}}\leq s-$ 1.
Equivalently, letting
$R^{r}\stackrel{{\scriptstyle\varphi}}{{\longrightarrow}}R^{m}\longrightarrow E\rightarrow 0$
stand for a finite free presentation of $E$ over $R$, $E$ satisfies condition $G_{s}$ if
$\operatorname{ht}I_{j}(\varphi)\geq m-j-(e-2),\;\text{\rm for}\;m-s-(e-2)\leq j\leq m-e,$ (3)
where $\operatorname{ht}\mathcal{I}$ denotes the height of the ideal $\mathcal{I}$.
The definition as applied to an ideal with positive grade translates, respectively, into any of the requirements
$\mu(I_{\mathfrak{p}})\leq\operatorname{ht}{\mathfrak{p}},\;\text{\rm for}\;{\mathfrak{p}}\in\operatorname{Spec}(R)\;\text{\rm with}\;1\leq\operatorname{ht}{\mathfrak{p}}\leq s-1.$ (4)
or
$\operatorname{ht}I_{j}(\varphi)\geq m-j+1,\;\text{\rm for}\;m-s+1\leq j\leq m-1.$ (5)
Moreover, in this regard, the following corollary of [16, Theorem 5.1] comes up quite often.
Proposition 2.2.
Let $R$ be a standard graded Cohen–Macaulay ring over a field and let $I\subset R$ denote an equigenerated perfect homogeneous ideal of codimension $2$ If $I$ satisfies condition $G_{s}$ then ${\rm indeg}(Q)\geq s.$
Recall that, over a Cohen–Macaulay base ring “codimension” and “height” are used interchangeably.
In particular, if $\mu(I)>\dim R$ and $I$ satisfaz condition $G_{s}$, then $\mu(JI^{j})<\mu(I^{j+1})$ for all $20\leq j\leq s-$ 2, where $J$ denotes a minimal reduction of $I$. That is, the reduction number of $I$ is at least $s-1$.
Quite a bit has been said long ago about the dimension of $\mathcal{S}_{R}(E)$ in regard to the behavior of the ideals of minors of the matrix $\varphi$ (see [12], [13]). Next is a somewhat technical, but useful result, that does not seem to have been recorded before in this precision.
Proposition 2.3.
Let $R$ be a standard graded Cohen-Macaulay ring of dimension $d$ and let $E$ a be an $R$-module with rank $e$. Assume the following conditions::::
  1. (1)
    $E$ has projective dimension one, with minimal free resolution
    $0\to R^{n-e}\stackrel{{\scriptstyle\varphi}}{{\longrightarrow}}R^{n}\to E\to 0,$
    such that $n\geq d+e-1$.
  2. (2)
    $E$ satisfies conditin $G_{s},$ for some $11\leq s\leq d-$ 1 .
  3. (3)
    There is a prime ideal $I_{1}(\varphi)\subset\mathfrak{q}\subset R$ with $\operatorname{ht}\mathfrak{q}\geq s$ such that $\operatorname{ht}I_{j}(\varphi)\geq\operatorname{ht}\mathfrak{q}-j+1$ for every $1\leq j\leq\operatorname{ht}\mathfrak{q}-s$.
Then $\dim\mathcal{S}_{R}(E)=n+\dim R/\mathfrak{q}.$
Proof.
We apply the Huneke–Rossi formula ([6])
$\dim{\mathcal{S}}_{R}(E)={\rm sup}_{\mathfrak{p}\in{\rm Spec}R}\left\{\dim R/\mathfrak{p}+\mu(E_{\mathfrak{p}})\right\}.$
$\bullet$ $\operatorname{ht}\mathfrak{p}\leq s-1$
$\dim R/\mathfrak{p}+\mu(I_{\mathfrak{p}})$ $=$ $d-\operatorname{ht}\mathfrak{p}+\mu(E_{\mathfrak{p}})$ (6)
$\leq$ $d+e-1\quad\quad\quad\quad\quad\quad\quad(\mbox{because $E$ satisfies $G_{s}$})$
$\leq$ $n\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\,\,(\mbox{byhypothesis}).$
$\bullet$ $s\leq\operatorname{ht}\mathfrak{p}<\operatorname{ht}\mathfrak{q}.$
Set $j:=\operatorname{ht}\mathfrak{q}-\operatorname{ht}\mathfrak{p}.$ In particular, $1\leq j\leq\operatorname{ht}\mathfrak{q}-s.$ So, by hypothesis we have $\operatorname{ht}I_{j}(\varphi)\geq\operatorname{ht}\mathfrak{q}-j+1=\operatorname{ht}\mathfrak{p}+1.$ From this, $I_{j}(\varphi)\not\subset\mathfrak{p}.$ Thus, $\mu(E_{\mathfrak{p}})\leq n-j=n-\operatorname{ht}\mathfrak{q}+\operatorname{ht}\mathfrak{p}.$ Therefore,
$\dim R/\mathfrak{p}+\mu(E_{p})\leq d-\operatorname{ht}\mathfrak{p}+n-\operatorname{ht}\mathfrak{q}+\operatorname{ht}\mathfrak{p}=n+d-\operatorname{ht}\mathfrak{q}=n+\dim R/\mathfrak{q}$ (7)
$\bullet$ $\mathfrak{p}=\mathfrak{q}$
Since $I_{1}(\varphi)\subset\mathfrak{q}$ then
$0\to R_{\mathfrak{q}}^{n-e}\stackrel{{\scriptstyle\varphi}}{{\longrightarrow}}R_{\mathfrak{q}}^{n}\to E_{\mathfrak{q}}\to 0$
is a minimal free resolution of $E_{\mathfrak{q}}$ over $R_{\mathfrak{p}}.$ Therefore, $\mu(E_{\mathfrak{q}})=n$ and
$\dim R/\mathfrak{q}+\mu(E_{\mathfrak{q}})=n+\dim R/\mathfrak{q}.$
$\bullet$ $\operatorname{ht}\mathfrak{p}\geq\operatorname{ht}\mathfrak{q}$
In this case we have $\dim R/\mathfrak{p}\leq\dim R/\mathfrak{q}$ and $\mu(E_{\mathfrak{p}})\leq n.$ Thus,
$\dim R/\mathfrak{p}+\mu(E_{\mathfrak{p}})\leq n+\dim R/\mathfrak{q}$
Having gone through all possibilities, we have $\dim\mathcal{S}_{R}(E)=n+\dim R/\mathfrak{q}$ as claimed. ∎
3 Preliminaries
We keep the assumptions and notation of the previous section, where $R=k[x_{1},\ldots,x_{d}]$ is a standard graded polynomial ring and $I\subset R$ is a perfect ideal of codimension $2$ with an $n\times(n-1)$ linear presentation matrix $\varphi$. In addition, we stipulate that throughout $I_{1}(\varphi)=\langle x_{1},\ldots,x_{d}\rangle$ and the ideal $I$ satisfies condition $G_{d-1}$, but not condition $G_{d}$.
Next is a first elementary output of this landscape.
Lemma 3.1.
With these assumptions, one has::::
  1. (1)
    $\operatorname{ht}I_{n-d+1}(\varphi)=d-1$.
  2. (2)
    There is a least integer $1\leq u\leq n-d$ such that $\operatorname{ht}I_{u+1}(\varphi)=d-1$. Thus, $\operatorname{ht}I_{i}(\varphi)=d$ for every $i\leq u$.
  3. (3)
    For a minimal prime $\mathfrak{p}$ of $R/I_{u+1}(\varphi)$, the rank of $\varphi$ over $R/\mathfrak{p}$ is $u$.
  4. (4)
    $\dim\mathcal{S}_{R}(I)=n$ and the Jacobian dual matrix $B$ of $\varphi$ has rank $d$ over $k[{\bf t}]$.
Proof.
(1) Since $I$ satisfies $G_{d-1}$ then, by (5), one has $\operatorname{ht}I_{n-d+2}(\varphi)\geq n-(n-d+2)+1=d-1$. Since $I_{n-d+2}(\varphi)\subset I_{n-d+1}(\varphi)$, then $\operatorname{ht}I_{n-d+1}(\varphi)\geq d-1$ as well. On the other hand, if $G_{d}$ fails at some $n-d+2\leq j\leq n-1$ then $G_{d-1}$ also fails at this value. Therefore, $G_{d}$ must fail at the value $j=n-d+1$, that is, $\operatorname{ht}I_{n-d+1}(\varphi)<n-(n-d+1)+1=d$.
(2) This is an immediate consequence of the first item.
(3) Let $r$ denote the rank of $\varphi$ over $R/\mathfrak{p}$, that is, $r=\min\{t\,|\,I_{t}(\varphi)\not\subset\mathfrak{p}\}.$ Since $I_{u+1}(\varphi)\subset\mathfrak{p}$ and $I_{u}(\varphi)\not\subset\mathfrak{p}$ (because $\operatorname{ht}I_{u}(\varphi)=d$) the statement follows. (Note en passant that the role of $u$ resembles the notion of the chaos invariant of [4].)
(4) Picking $\mathfrak{q}=\langle x_{1},\ldots,x_{d}\rangle$, we have
  1. $\bullet$
    $d+1-1<n$
  2. $\bullet$
    $I_{1}(\varphi)=\mathfrak{q};$
  3. $\bullet$
    $\operatorname{ht}I_{j}(\varphi)\geq\operatorname{ht}\mathfrak{q}-j+1$ for every $11\leq j\leq\operatorname{ht}\mathfrak{q}-(d-1)=$ 1 .
Thus, by Proposition 2.3, $\dim\mathcal{S}_{R}(I)=n+\dim R/\mathfrak{q}=n.$
Now, to show that the rank of $B$ is $d$, let $M$ be the cokernel of $B.$ Thus, we have the finite presentation
$k[{\bf t}]^{n-1}\stackrel{{\scriptstyle B}}{{\longrightarrow}}k[{\bf t}]^{d}\to M\to 0.$
By the equality ${\bf t}\varphi={\bf x}B$ we deduce $\mathcal{S}_{R}(I)\simeq\mathcal{S}_{k[{\bf t}]}(M).$ Thus, $\dim\mathcal{S}_{k[{\bf t}]}(M)=n.$ Hence, once more by the Huneke-Rossi formula, localizing at the zero ideal:
$n+\mu(M_{(0)})\leq n.$
Therefore, ${\rm rk}(M)=\mu(M_{(0)})=0.$ Consequently, ${\rm rk}(B)=d-{\rm rk}(M)=d.$
Next is how reading off the above departure let us set $\varphi$ in a more convenient form (see [4, The proof of Theorem 2.4]).
Lemma 3.2.
($k$ algebraically closed) The matrix $\varphi$ is conjugate to a matrix of the following form
[…]
where the $a_{i}$’s and the blank entries are linear forms in $k[x_{2},\dots,x_{d}]$.
Proof.
Let $\mathfrak{p}$ be a minimal prime of $R/I_{u+1}(\varphi)$. Since $k$ is algebraically closed and $\operatorname{ht}I_{u+1}(\varphi)=d-1$, by a change of variables we may assume that $\mathfrak{p}=\langle x_{2},\ldots,x_{d}\rangle$.
Write $\varphi=x_{1}\varphi_{1}+\cdots+x_{d}\varphi_{d},$ where $\varphi_{1},\ldots,\varphi_{d}$ are $n\times(n-1)$ matrices over $k.$ Note that $x_{1}\varphi_{1}$ is same as $\varphi$ taken over $R/\mathfrak{p}.$ Thus, $u=\operatorname{rank}\varphi_{1}$ by Lemma 3.1 (3). Hence, $\varphi_{1}$ is conjugate to
$\left[\begin{array}[]{cccc}\mathbb{I}_{u}&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{0}\end{array}\right],$
by means of matrices $A\in{\rm GL}_{n}(k)$ on the left, and $B\in{\rm GL}_{n-1}(k)$ on the right. Then $\varphi$ is conjugate by these same matrices to one of the desired form. ∎
We assume throughout the rest of the paper that $\varphi$ is as in (8).
4 The proof of Proposition 1.3
We keep the notation and assumptions of Section 3, in which $I\subset R=k[x_{1},\ldots,x_{d}]$ was a height two linearly presented perfect ideal, satisfying $G_{d-1}$ but not $G_{d}$.
Since we are assuming that its Hilbert–Burch matrix $\varphi$ is as in (8), the corresponding Jacobian dual matrix $B$ acquires the following form:
$B=\left[\begin{array}[]{cccc}t_{1}&\cdots&t_{u}&{\bf 0}\\ &&&B^{\prime}\end{array}\right],$ (9)
where $B^{\prime}$ is a $(d-1)\times(n-1-u)$ matrix, and the blank entries are linear forms in $k[{\bf t}]$.
Let $\varphi_{u}$ denote the submatrix of $\varphi$ omitting the first $u$ columns, giving rise to a short exact sequence of $R$-modules
$R^{n-1-u}\stackrel{{\scriptstyle\varphi_{u}}}{{\longrightarrow}}R^{n}\to E\to 0.$ (10)
Next we introduce the main properties of the $R$-module $E$, the Jacobian dual matrix $B$ and its submatrix $B^{\prime}$.
Lemma 4.1.
Let $E$ be as in (10). Then:
  1. (1)
    $E$ has rank $u+1$ (i.e., $E$ has projective dimension one).
  2. (2)
    $E$ satisfies the $G_{d-1}$ condition.
  3. (3)
    $\dim{\mathcal{S}}_{R}(E)=n+1.$
Proof.
(1) Let $j\geq 1$ be an integer such that $j+u\leq n-1$. Any $(j+u)$-minor of $\varphi$ involves at least $j$ columns of $\varphi_{u}.$ Thus, by Laplace, every $(j+u)$-minor of $\varphi$ is a linear combination over $R$ of $j$-minors of $\varphi_{u},$ that is, $I_{j+u}(\varphi)\subset I_{j}(\varphi_{u}).$
On the other hand, suppose in addition that $n-(u+1)-(d-1)+2\leq j\leq n-(u+1),$ i.e., $n-(d-1)+1\leq j+u\leq n-1$. Since $I$ satisfies condition $G_{d-1}$, then ${\rm ht\,}I_{j+u}(\varphi)\geq n-(j+u)+1.$ Thus, ${\rm ht\,}I_{j}(\varphi_{u})\geq n-(j+u)+1$ as well. In particular, for $j=n-u-1$, we find that $\operatorname{ht}I_{n-u-1}(\varphi_{u})\geq 2.$ This implies that ${\rm rk\,}E=u+1$.
(2) By the argument in the previous item, now applying the criterion of (3), with ${\rm rk\,}E=u+1$, we have that $E$ satisfies the $G_{d-1}$ condition.
(3) By the previous item, $E$ is an $R$-module with rank $e=u+1$ having minimal graded free resolution
$0\to R^{n-1-u}\stackrel{{\scriptstyle\varphi_{u}}}{{\longrightarrow}}R^{n}\to E\to 0$
satisfying $G_{d-1}.$ Note that
  1. $\bullet$
    $d+e-1\leq n$ because $u\leq n-d$ and $e=u+1;$
  2. $\bullet$
    $I_{1}(\varphi_{u})\subset\mathfrak{q}:=\langle x_{2},\ldots,x_{d}\rangle;$
  3. $\bullet$
    $\operatorname{ht}I_{j}(\varphi_{u})\geq\operatorname{ht}\mathfrak{q}-j+1$ for every $1\leq j\leq\operatorname{ht}\mathfrak{q}-(d-1)$ (this follows by emptiness).
Then, by Proposition 2.3 with $\mathfrak{q}:=\langle x_{2},\ldots,x_{d}\rangle,$ we have $\dim\mathcal{S}(E)=n+\dim R/\mathfrak{q}=n+1.$
In addition, one has:
Lemma 4.2.
The $(d-1)\times(n-1-u)$ matrix submatrix $B^{\prime}$ in (9) has maximal rank.
Proof.
Set $A:=\left[\begin{matrix}\boldsymbol{0}\\ B^{\prime}\end{matrix}\right]$ for lighter reading, where $\boldsymbol{0}$ is the $1\times(n-u-1)$ null matrix, and let $M$ denote the cokernel of the map of $k[{\bf t}]$-modules defined by $A$, namely:
$k[{\bf t}]^{n-u-1}\stackrel{{\scriptstyle A}}{{\longrightarrow}}k[{\bf t}]^{d}\to M\to 0.$
Clearly, ${\rm rk\,}M=d-{\rm rk}\,A=d-{\rm rk}\,B^{\prime}\geq 1.$ On the other hand, $A$ is a Jacobian dual like matrix relative to $\varphi_{u}$, that is,
${\bf t}\varphi_{u}={\bf x}A.$
Therefore, ${\mathcal{S}}_{k[{\bf t}]}(M)\simeq{\mathcal{S}}_{R}(E).$ Thus, $\dim{\mathcal{S}}_{k[{\bf t}]}(M)=n+1.$ Hence, once more by by the Huneke-Rossi formula, localizing at the zero ideal:
$n+\mu(M_{(0)})\leq n+1,$
that is, ${\rm rk\,}M={\rm rk}\,M_{(0)}=\mu(M_{(0)})\leq 1.$ Therefore, ${\rm rk\,}M)=1.$ In particular, ${\rm rk\,}B^{\prime}=d-1.$
We now prove Proposition 1.3, restated for convenience.
Proposition 4.3.
Let $Q$ denote the homogeneous defining ideal of the especial fiber $\mathcal{F}_{R}(I)$. Then the rational map $\psi:\mathbb{P}^{d-1}\dasharrow\mathbb{P}^{n-1}$ defined by the linear span of $I$ is birational onto the image.
Proof.
Note that the image of $\psi$ is identified with $\mathcal{F}_{R}(I)$. The proof follows pretty much the steps of the proof of [4, Theorem 2.4], via the birationality criterion of [3, Theorem 2.18]. Let then $Q\subset k[{\bf t}]$ denote the homogeneous defining ideal of $\mathcal{F}_{R}(I)$. Namely, we will show that there is a square submatrix $M$ of the matrix $B$ in (9), of order $d-1$, such that $\det(M)\neq 0\text{ (mod }Q)$, that is, ${\operatorname{rank}\,}_{\mathcal{F}(I)}(B)\geq d-1$. From this, the result follows by the above criterion.
By Lemma 4.2, $I_{d-1}(B^{\prime})\neq 0.$ Thus, since $I_{d-1}(B^{\prime})\subseteq I_{d-2}(B^{\prime}),$ there is a square submatrix $M^{\prime}$ of $B^{\prime}$ of order $d-2$ such that $\det(M^{\prime})\neq 0.$ Let
$M=\left[\begin{array}[]{cc}t_{n}&0\\ &M^{\prime}\end{array}\right],$
a square submatrix of $B$ of order $d-1$. Clearly, $\det(M)=t_{n}\det(M^{\prime})\neq 0$.
We assert that $\det(M)\not\subset Q$. Otherwise, since $Q$ is a prime, then either $t_{u}\in Q$ or else $\det(M^{\prime})\in Q$, both of degrees strictly less than $d-1$. But, since $I$ satisfies $G_{d-1}$, Proposition 2.2 says that the initial degree of $Q$ is at least $d-1$.
[…]
Corollary 4.4.
The following statements hold::::
  1. (i)
    There is a $d\times(d-1)$ submatrix $\mathcal{B}$ of $B$ of rank $d-1$ modulo $Q$.
  2. (ii)
    For every $d\times(d-1)$ submatrix $\mathcal{B}$ of $B$ of rank $d-1$ modulo $Q$, the sequence of its ordered signed $(d-1)$-minors gives a representative of the inverse of $\psi.$
  3. (iii)
    Let $\mathcal{B}$ denote any matrix as in the previous item, and let ${\Delta}$ stand for the sequence of its the ordered signed $(d-1)$-minors. Then, regarding $\mathcal{R}_{R}(I)$ as an $R[{\bf t}]/Q$-algebra, one has an isomorphism of $R[{\bf t}]/Q$-algebras
    $\partial_{\mathcal{B}}:\mathcal{R}_{R}(I){\stackrel{{\scriptstyle\sim}}{{\rightarrow}}}\mathcal{R}_{k[{\bf t}]/Q}(\langle{\Delta},Q\rangle/\langle Q\rangle).$ (11)
Proof.
(i) is essentially an outcome of the argument in the proof of (4.3), while (ii) stems from [3, Theorem 2.18, (a) and Supplement (ii)].
(iii) is [14, Proposition 2.1]. Note, for further clarification, that $\mathcal{R}_{R}(I)\simeq R[{\bf t}]/\mathcal{J}$, where $\mathcal{J}$ contains $Q$, so $\mathcal{R}_{R}(I)$ is naturally an $R[{\bf t}]/Q$-algebra. ∎
Remark 4.5.
(1) In item (iii) of the above corollary the ordered sequence of $(d-1)$-minors of $\Delta$ is a representative of the inverse map to the rational map in the main proposition.
(2) We likewise emphasize an important aspect of the isomorphism $\partial_{\mathcal{B}}$ in (11) , where its restriction $\widetilde{\partial_{\mathcal{B}}}$ to $R=k[x_{1},\dots,x_{d}]$ entails the relations ~$\widetilde{\partial_{\mathcal{B}}}(x_{1})=\delta_{1}(\bmod\;Q),\dots,\widetilde{\partial_{\mathcal{B}}}(x_{d})=\delta_{d}(\bmod\;Q)$. This understood, letting ${\bf\Delta}$ denote the residue class of $\Delta$ modulo $Q$, we take liberty in writing it as ${\bf x}\mapsto{\bf\Delta}$.
5 Proof of the main result
A tall side of this section is a specialization toward one dimension less.
Since $\operatorname{ht}I_{n-d+1}(\varphi)=d-1$ (Lemma 3.1(1)) and $I_{n-d+1}(\varphi)\supseteq I_{n-d+2}(\varphi)\supseteq\cdots\supseteq I_{n-1}(\varphi)=I$ we have $\operatorname{ht}I_{j}(\varphi)\leq d-1$ for $n-d+2\leq j\leq n-1.$ Thus, by the prime avoidance, there is linear form $\mathfrak{f}=x_{1}-b_{2}x_{2}-\cdots-b_{d}x_{d}\in k[x_{1},\dots,x_{d}]_{1}$ avoiding the minimal primes of the minors ideal $I_{j}(\varphi)$ for $n-d+2\leq j\leq n-1.$ In particular, $\mathfrak{f}$ is regular over $R/I$ and $\dim R/\langle I_{j}(\varphi),\mathfrak{f}\rangle=\dim R/I_{j}(\varphi)-1$ for every $n-d+2\leq j\leq n-1.$ In particular, the natural projection $R\twoheadrightarrow R/\langle\mathfrak{f}\rangle\simeq k[x_{2},\dots,x_{d}]=:\overline{R}$ maps $I$ onto the height $2$ perfect ideal $\overline{I}\subset\overline{R}$ satisfying $G_{d-1}$, with an $n\times(n-1)$ syzygy (Hilbert–Burch) matrix over $k[x_{2},\dots,x_{d}]$ of the following form
$\overline{\varphi}=\left[\begin{array}[]{cccccc}b_{2}x_{2}+\cdots+b_{d}x_{d}+a_{1}&&&&&\\ &\ddots&&&&\\ &&b_{2}x_{2}+\cdots+b_{d}x_{d}+a_{u}&&&\\ &&&&&\\ \end{array}\right],$ (12)
where the $a_{i}$’s and the blank entries are linear forms. Thus, by a similar token, the corresponding Jacobian dual matrix – namely, the unique $(d-1)\times(n-1)$ matrix $\overline{B}$ with linear entries in $k[{\bf t}]$ such that ${\bf t}\cdot\overline{\varphi}=[x_{2}\cdots x_{d}]\overline{B}$ – has the following format:
$\overline{B}=\left[\begin{array}[]{cccccc}\ell_{21}+b_{2}t_{1}&\cdots&\ell_{2u}+b_{2}t_{u}&\ell_{2,u+1}&\cdots&\ell_{2,n-1}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ \ell_{d1}+b_{d}t_{1}&\cdots&\ell_{du}+b_{d}t_{u}&\ell_{d,u+1}&\cdots&\ell_{d,n-1}\end{array}\right],$ (13)
where $\ell_{i,j}$ is a linear form in $k[{\bf t}]$.
Now, given a $d\times(d-1)$ submatrix $\mathcal{B}$ as in Corollary 4.4 (i), having rank $d-1$ modulo $Q$, let $\overline{\mathcal{B}}$ stand for the “corresponding” $(d-1)\times(d-1)$ submatrix of $\overline{B}$. That is to say, if $\mathcal{B}$ is based on the columns of $B$ indexed by ${\bf v}=v_{1},\dots,v_{d-1}$ ($11\leq v_{1}<\cdots<v_{d-1}\leq n-$ 1), then $\overline{\mathcal{B}}$ is based on the columns of $\overline{B}$ indexed similarly.
Lemma 5.1.
With the above notation along with Proposition 4.3, one has::::
  1. (i)
    The ideal $\langle I_{1}({\bf t}\cdot\overline{\varphi}),I_{d-1}(\overline{B})\rangle\subset\overline{R}[{\bf t}]$ is a bihomogeneous defining ideal of the Rees algebra of $\overline{I}$ over $\overline{R}$.
  2. (ii)
    For every $d\times(d-1)$ submatrix $\mathcal{B}$, one has $\det(\overline{\mathcal{B}})=i^{-1}(\widetilde{\partial_{\mathcal{B}}}(\mathfrak{f}))$, the pull back via the canonical surjection $i:k[{\bf t}]\twoheadrightarrow k[{\bf t}]/Q$.
Proof.
(i) The first part is due to [11, Theorem 1.3], since $\overline{I}$ satisfies $G_{d-1}$ and $k$ is infinite.
(ii) Say, $\mathcal{B}=\left[\begin{matrix}\ell_{1v_{1}}&\cdots&\ell_{dv_{1}}\\ \vdots&\ddots&\vdots\\ \ell_{1v_{d-1}}&\cdots&\ell_{dv_{d-1}}\end{matrix}\right]^{\rm tr}$, for certain $11\leq v_{1}<\cdots<v_{d-1}\leq n-$ 1. Then,
$\det(\overline{\mathcal{B}})$ $=$ $\det\left[\begin{matrix}\ell_{2v_{1}}+b_{2}\ell_{1v_{1}}&\cdots&\ell_{2v_{d-1}}+b_{2}\ell_{1v_{d-1}}\\ \vdots&\ddots&\vdots\\ \ell_{dv_{1}}+b_{d}\ell_{1v_{1}}&\cdots&\ell_{dv_{d-1}}+b_{d}\ell_{1v_{d-1}}\\ \end{matrix}\right]=\det\left[\begin{matrix}\ell_{2v_{1}}&\cdots&\ell_{2v_{d-1}}\\ \vdots&\ddots&\vdots\\ \ell_{dv_{1}}&\cdots&\ell_{dv_{d-1}}\end{matrix}\right]$
$+b_{2}\det\left[\begin{matrix}\ell_{1v_{1}}&\cdots&\ell_{1v_{d-1}}\\ \ell_{3v_{1}}&\cdots&\ell_{3v_{d-1}}\\ \vdots&\ddots&\vdots\\ \ell_{dv_{1}}&\cdots&\ell_{dv_{d-1}}\end{matrix}\right]+\cdots+b_{d}\det\left[\begin{matrix}\ell_{2v_{1}}&\cdots&\ell_{2v_{d-1}}\\ \vdots&\ddots&\vdots\\ \ell_{d-1,v_{1}}&\cdots&\ell_{d-1,v_{d-1}}\\ \ell_{1v_{1}}&\cdots&\ell_{1v_{d-1}}\end{matrix}\right]=i^{-1}(\widetilde{\partial_{\mathcal{B}}}(\mathfrak{f})),$
the result of applying the lifting $\widetilde{\partial_{\mathcal{B}}}$ of the isomorphism (11) to the linear form $\mathfrak{f}\in R$ and pulling back to $k[{\bf t}]$. Notice that, $\ell_{1j}=t_{j}$ for $1\leq j\leq u$ and $\ell_{1j}=0$ otherwise. ∎
We state again the main result.
Theorem 5.2.
With the notation established so far, let $\mathcal{J}$ and $Q$ respectively denote the bihomogeneous defining ideal of the Rees algebra $\mathcal{R}_{R}(I)$ and the homogeneous defining ideal of the special fiber $\mathcal{F}_{R}(I)$, and let $B$ denote the associated Jacobian dual matrix. Then::::
  1. (i)
    $\mathcal{J}=\langle I_{1}({\bf t}\cdot\varphi),Q\rangle$ as ideals of $R[{\bf t}]$, i.e., $I$ is an ideal of fiber type.
  2. (ii)
    The initial degree of $Q$ is $d-1;$ in particular, the inclusion of ideals $\langle I_{1}({\bf t}\cdot\varphi),I_{d}(B)\rangle\subset\langle I_{1}({\bf t}\cdot\varphi),Q\rangle$ is proper. Hence, the Rees algebra of $I$ is not of the expected type.
Proof.
As discussed priorly, the goal is to prove that $\mathcal{J}\subseteq\langle I_{1}({\bf t}\cdot\varphi),Q\rangle$. The strategy will consist in induction on the sum $p+q$, where $(p,q)$ is the bidegree of a bihomogeneous element $\mathbf{f}\in\mathcal{J}$. Our starting step is the value $p+q\leq 2$ as for these values $\mathbf{f}\in\langle I_{1}({\bf t}\cdot\varphi),Q\rangle$.
Suppose then that $p+q\geq 3$. Consider the canonical commutative diagram below
$\begin{array}[]{ccccccc}0\to&\mathcal{J}&\to&R[{\bf t}]&\to&\mathcal{R}(I)&\to 0\\ &\downarrow&&\downarrow&&\downarrow&\\ 0\to&\overline{\mathcal{J}}&\to&\overline{R}[{\bf t}]&\to&\mathcal{R}_{\overline{R}}(\overline{I})&\to 0\end{array},$
where the downarrows are surjetive. Applying Lemma 5.1 (i) gives $\overline{\mathbf{f}}\in\langle I_{1}({\bf t}\cdot\overline{\varphi}),I_{d-1}(\overline{B})\rangle\subset\overline{\mathcal{J}}$. Say, $\overline{\mathbf{f}}=\overline{\mathbf{g}}+\overline{\mathbf{h}}$, where $\overline{\mathbf{g}}\in I_{1}({\bf t}\cdot\overline{\varphi})$ and $\overline{\mathbf{h}}\in I_{d-1}(\overline{B})$.
Now, by Lemma 5.1 (ii), we can write
$\mathbf{h}$ $=$ $\mathfrak{q}+\sum_{i=1}^{r}\mathfrak{c}_{i}\det\overline{\mathcal{B}_{i}}$
$=$ $\mathfrak{q}+\sum_{i=1}^{r}\mathfrak{c}_{i}\;i^{-1}(\widetilde{\partial_{\mathcal{B}_{i}}}(\mathfrak{f})),$
for certain $\mathfrak{c}_{i}\in R[{\bf t}]$, and where $\mathfrak{q}\in Q,$ and $\{\mathcal{B}_{1},\ldots,\mathcal{B}_{r}\}$ is the total set of matrices as in Proposition 4.3 (iii).
Following the last line in Remark 4.5 (2), write $i^{-1}(\widetilde{\partial_{\mathcal{B}_{i}}}(\mathfrak{f}))=\mathfrak{f}({\Delta}_{i})$ for simplicity. Since minors commute with homomorphisms, lifting back to $R[{\bf t}]$ through the above diagram, one has
$\mathbf{f}-\mathbf{g}=\mathfrak{q}+\sum_{i=1}^{r}{\mathfrak{c}_{i}}\,\mathfrak{f}({\Delta}_{i})+\mathfrak{p}\mathfrak{f}$ (14)
for a suitable bihomogeneous form $\mathfrak{p}\in R[{\bf t}].$
Since both $\mathbf{g}$ and $\mathfrak{q}$ belong to $\langle I_{1}({\bf t}\cdot\varphi),Q\rangle$, it is enough to show that $\mathbf{f}_{1}:=\sum_{i=1}^{r}{\mathfrak{c}_{i}}\,\mathfrak{f}({\Delta}_{i})+\mathfrak{p}\mathfrak{f}\in\langle I_{1}({\bf t}\cdot\varphi),Q\rangle$.
As $\mathbf{f}$ is bihomogeneous of bidegree $(p,q)$, so the summands have to be of bidegree $(p,q)$ as well, hence $\mathfrak{c}_{1},\ldots,\mathfrak{c}_{r}$ are bihomogeneous of bidegree $(p,q-2)$ since a minor of $\Delta_{i}$ has degree $d-1$ in ${\bf t}$, and $\mathfrak{p}$ are bihomogeneous of bidegree $(p-1,q).$
We next take $i=1$ and evaluate ${\bf x}\mapsto\Delta_{1}$, to get
$\mathbf{f}_{1}(\Delta_{1},{\bf t})=\sum_{i=1}^{r}{\mathfrak{c}_{i}}(\Delta_{1},{\bf t})\,\mathfrak{f}({\Delta}_{i})+\mathfrak{p}(\Delta_{1},{\bf t})\mathfrak{f}(\Delta_{1}),$
and back by restriction of the inverse homomorphism (11):
$\underbrace{\sum_{i=2}^{r}\mathfrak{c}_{i}\,\mathfrak{f}({\Delta}_{i})+\mathfrak{p}\mathfrak{f}}_{:=\mathbf{f}_{2}}+\mathfrak{c}_{1}\mathfrak{f}\in\mathcal{J}.$
Since $\mathcal{J}$ is a bihomogeneous ideal, and $\mathbf{f}_{2}$ and $\mathfrak{c}_{1}\mathfrak{f}$ have bidegrees $(p,q)$ and $(p+1,q-2)$, respectively, then both are elements of $\mathcal{J}.$ In addition, $\mathfrak{f}\notin\mathcal{J},$ hence $\mathfrak{c}_{1}\in\mathcal{J}.$ But $p+q-2<p+q$, hence $\mathfrak{c}_{1}\in\langle I_{1}({\bf t}\cdot\varphi),Q\rangle$ by the inductive hypothesis. Now, using the fact that $\mathbf{f}_{2}\in\mathcal{J}$ we can use the evaluation ${\bf x}\mapsto{\bf g}_{2}$ to conclude as before that $\mathfrak{c}_{2}\in\langle I_{1}({\bf t}\cdot\varphi),Q\rangle.$ In this manner we eventually obtain that $\{\mathfrak{c}_{1},\ldots,\mathfrak{c}_{r}\}\subset\langle I_{1}({\bf t}\cdot\varphi),Q\rangle$, hence $\mathfrak{p}\in\langle I_{1}({\bf t}\cdot\varphi),Q\rangle$ as well, as was to be shown.
(ii) Consider the following $d\times(n-u)$ submatrix of $B$
$\tilde{B}=\left[\begin{matrix}t_{u}&\boldsymbol{0}\\ \ast&B^{\prime}\end{matrix}\right].$
By Lemma 4.2, $I_{d-1}(B^{\prime})\neq 0$. Therefore, $I_{d}(\tilde{B})=t_{u}I_{d-1}(B^{\prime})\neq 0$ as well. But, $I_{d}(\tilde{B})\subset I_{d}(B)$, while the latter is contained in the bihomogeneous defining ideal of $\mathcal{R}_{R}(I)$. Therefore, by item (i), one must have an inclusion $I_{d}(B)\subset Q$. Since $Q$ is a prime ideal generated in degree $\geq d-1$, we must conclude that $I_{d-1}(B^{\prime})\subset Q$. This shows that the initial degree of $Q$ is $d-1$. In particular, as $I_{d}(B)$ lives in degree $d$, the inclusion $\langle I_{1}({\bf t}\cdot\varphi),I_{d}(B)\rangle\subset\langle I_{1}({\bf t}\cdot\varphi),Q\rangle$ is proper. ∎